OpenStudy (anonymous):

Are there any values of t so that sin t= 5/4?

ganeshie8 (ganeshie8):

hint : \(-1 \le \sin(t) \le 1\)

OpenStudy (anonymous):

So there isn't any number that t could equal?

ganeshie8 (ganeshie8):

you're right, there is no value that ever makes sin t greater than 1

OpenStudy (anonymous):

Okay, thank you for the explanation! :D

ganeshie8 (ganeshie8):

np.. .u wlc :)